OpenStudy (anonymous):

Find the limit of sin(ax)/tan(bx) as x approaches 0.

OpenStudy (anonymous):

a/b

OpenStudy (anonymous):

thanks, but what's the solution?

OpenStudy (anonymous):

sin(ax)/tan(b(x)=(sin(ax)cos(b(x))/sin(bx) then l'hopital,lim=lim(acos(ax)cos(bx)-bsin(ax)sin(b(x))/bcos(bx)=1 since the sines are 0

OpenStudy (anonymous):

sorry =a/b

OpenStudy (anonymous):

thanks. we still haven't reached this part in class but i'll try to understans and learn. :)

OpenStudy (anonymous):

understand*

OpenStudy (anonymous):

good luck:)

OpenStudy (zarkon):

\[\frac{\sin(ax)}{\tan(bx)}=\frac{\sin(ax)}{\sin(bx)/\cos(bx)}=\frac{\sin(ax)}{1}\frac{1}{\sin(bx)}\cos(bx)\] \[=a\frac{\sin(ax)}{ax}\frac{1}{b}\frac{bx}{\sin(bx)}\cos(bx)\] letting \[x\to 0\] \[=a\cdot1\cdot\frac{1}{b}\cdot1\cdot1=\frac{a}{b}\]